# Designer three-dimensional electronic bands in asymmetric transition metal dichalcogenide heterostructures

### v1

@@
[…]
figurec
Designer three-dimensional electronic bands in asymmetric transition metal dichalcogenide heterostructures
Oliver J. Clark oliver.clark1@monash.edu School of Physics and Astronomy, Monash University, Clayton, VIC, Australia    Anugrah Azhar Department of Physics and Astronomy, University of Manchester, Oxford Road, Manchester M13 9PY, UK Physics Study Program, Faculty of Science and Technology, Syarif Hidayatullah State Islamic University Jakarta, Tangerang Selatan 15412, Indonesia    Ben A. Chambers Flinders Microscopy and Microanalysis, Flinders University, Bedford Park, South Australia 5042, Australia    Daniel McEwen School of Physics and Astronomy, Monash University, Clayton, VIC, Australia ARC Centre for Future Low Energy Electronics Technologies, Monash University, Clayton, VIC, Australia    Thi-Hai-Yen Vu School of Physics and Astronomy, Monash University, Clayton, VIC, Australia    M. Tofajjol H. Bhuiyan School of Physics and Astronomy, Monash University, Clayton, VIC, Australia    Rodion V. Belosludov Institute for Materials Research, Tohoku University, Sendai, 980-8577, Japan    Aaron Bostwick Advanced Light Source, Lawrence Berkeley National Laboratory, Berkeley, CA, 94720 USA    Chris Jozwiak Advanced Light Source, Lawrence Berkeley National Laboratory, Berkeley, CA, 94720 USA    Eli Rotenberg Advanced Light Source, Lawrence Berkeley National Laboratory, Berkeley, CA, 94720 USA    Seng Huat Lee Department of Physics, Pennsylvania State University, University Park, PA, 16802, USA 8 2D Crystal Consortium, Materials Research Institute, Pennsylvania State University, University Park, PA, 16802, USA    Zhiqiang Mao Department of Physics, Pennsylvania State University, University Park, PA, 16802, USA 8 2D Crystal Consortium, Materials Research Institute, Pennsylvania State University, University Park, PA, 16802, USA    Geetha Balakrishnan Department of Physics, University of Warwick, Coventry CV4 7AL, United Kingdom    Federico Mazzola Department of Molecular Sciences and Nanosystems, Ca Foscari University of Venice, Venice IT-30172, Italy CNR-SPIN UOS Napoli, Complesso Universitario di Monte Sant’Angelo, Via Cinthia 80126, Napoli, Italy    Sarah L. Harmer Institute for Nanoscale Science and Technology, Flinders University, Bedford Park, South Australia 5042, Australia Flinders Microscopy and Microanalysis, Flinders University, Bedford Park, South Australia 5042, Australia    Michael S. Fuhrer School of Physics and Astronomy, Monash University, Clayton, VIC, Australia ARC Centre for Future Low Energy Electronics Technologies, Monash University, Clayton, VIC, Australia    M. Saeed Bahramy Department of Physics and Astronomy, University of Manchester, Oxford Road, Manchester M13 9PY, UK    Mark T. Edmonds School of Physics and Astronomy, Monash University, Clayton, VIC, Australia ARC Centre for Future Low Energy Electronics Technologies, Monash University, Clayton, VIC, Australia ANFF-VIC Technology Fellow, Melbourne Centre for Nanofabrication, Victorian Node of the Australian National Fabrication Facility, Clayton, VIC 3168, Australia
@@
Van der Waals materials enable the construction of atomically sharp interfaces between compounds with distinct crystal and electronic properties Pham et al. (2022). This is dramatically exploited in moiré systems, where a lattice mismatch or twist between monolayers generates an emergent in-plane periodicity, giving rise to electronic properties absent in the constituent materials Devakul et al. (2021); Guo et al. (2023). In contrast, vertical superlattices—formed by stacking dissimilar van der Waals materials in the out-of-plane direction on the nanometer scale—remain largely unexplored, despite their potential to realize analogous emergent phenomena in three dimensions Wan et al. (2024). Here, we employ angle-resolved photoemission spectroscopy and density functional theory to investigate six-to-eight layer transition metal dichalcogenide (TMD) heterostructures. We find that even a single superlattice unit, i.e. a pair of stacked few-layer materials, exhibits emergent bands that closely resemble the hallmark three-dimensional electronic structure of many-layer TMDs, while simultaneously exhibiting an orbital- and momentum-dependent delocalization through the entire heterostructure due to its asymmetric architecture. We demonstrate how non-trivial band topology and unique TMD fermiologies are readily stabilized through these mechanisms, thus providing direct evidence for the novel electronic states possible through vertical superlattice engineering. These findings establish asymmetric heterostructures as platforms to invoke and control unprecedented combinations of the diverse quantum phases native to many-layer TMDs.
Figure 1: Chemical nature of an anisotropic transition metal heterostructure. (a) Schematic representation of a (WSe2)3/(MoSe2)3 heterostructure. Due to the $D_{3h}$ symmetry of each monolayer, a net dipole moment ±$\pm\mathbf{P}$ is induced, oriented perpendicular to the stacking axis, and oppositely aligned with respect to the adjacent layer. (b) and (c) Out-of-plane spin-projected electronic band structures for monolayers of WSe2 and MoSe2, respectively. The parameter $\Delta E_{\text{SO}}^{\text{W/Mo}}$ denotes the valley-dependent spin-orbit splitting at the K-point of the Brillouin zone. (d) and (e) Photoemission spectra collected from isolated 3ML thick flakes of WSe2 and MoSe2 (using 108 and 64 eV photons, respectively, and symmetrized about $\Gamma$) .The bands at $\Gamma$ are numbered from low energy to high. (f) and (g) Schematic energy diagrams illustrating the bonding and anti-bonding interactions at the K and $\Gamma$ points, respectively, with their orbital character predominantly influenced by the in-plane $d_{||}=\{{d_{xy},d_{x^{2}-y^{2}}}\}$ and out-of-plane $d_{z^{2}}$ orbitals of W/Mo. (h) Orbital-projected electronic structure of the (WSe2)3/(MoSe2)3 heterostructure, highlighting the distinct contributions from the constituent layers. The bands at $\Gamma$ are numbered from low energy to high.
@@
Transition metal dichalcogenides (TMDs) are popular components for functional devices due to the ease in which their vast array of intrinsic properties can be augmented Chhowalla et al. (2013); Ahn (2020); Ciarrocchi et al. (2022); Manzeli et al. (2017); Regan et al. (2022). This is facilitated through anisotropic bonding, enabling both the isolation of TMD flakes of arbitrary thicknesses and their subsequent recombination into stacks with atomically sharp interfaces without the need for epitaxial registry Lisi et al. (2021); Pei et al. (2022); Nunn et al. (2023); Stansbury et al. ; Jones et al. (2022). These procedures can both alter the electronic properties inherent to the component compounds, and induce otherwise-inaccessible correlated phenomena through the fine tuning of the laboratory-defined relative orientations, or ‘twist angle’ Devakul et al. (2021); Guo et al. (2023). Despite van der Waals bonding, however, the interlayer hopping along the $c$-axis in TMDs is significant, driving orbital-dependent bandwidths of up to several electron volts in bulk systems Clark et al. (2018, 2022); Riley et al. (2014); Watson et al. (2024). This is yet to be exploited in artificial heterostructures, with current research largely rooted to homo- and heterobilayer systems, and thus to TMD properties in the 2D limit. This precludes, for example, the fine-tuning of band gaps through band quantisation Splendiani et al. (2010); Mak et al. (2010); Sun et al. (2016), the induction of $k_{z}$-mediated topological phases Bahramy et al. (2018), or the modification of charge density wave and magnetic transitions dependent on interlayer coupling Zhang et al. (2025); Watson et al. (2019); Feng et al. (2018).
@@
Here, through nano-focused angle-resolved photoemission (nano-ARPES) and first principles calculations, we showcase vertical TMD heterostructures with emergent three-dimensional electronic band structures. Specifically, our 6-8 layer heterostructures, constructed simply by interfacing pairs of few-layer WSe2, MoSe2 and NbSe2 flakes, possess an orbital-dependent band localisation along the asymmetric $c$-axis. In this scenario, three-dimensional bands delocalised throughout the heterostructure coexist and interplay with layer-locked two-dimensional states to produce tunable arrays of material properties unlike those in either natural systems or artificial heterobilayers. We give special focus to $k_{z}$-mediated topological phase transitions, triggered by the interfacing of the few-layer TMD flakes, which produce pairs of distinct topologically non-trivial surfaces that are each defined largely by a single component material, unlike any naturally existing topological system. Similar phenomena should be expected for asymmetric heterostructures constructed from any pair of TMDs, and thus our findings unveil an enormous parameter space within which customizable three-dimensional electronic frameworks, mixing 2- and 3D properties from across the TMD family, can be produced and fine-tuned.
[…]
@@
Figure 1 overviews the mechanism underpinning our findings. Fig.1(a) illustrates the lattice structure of a 6-layer MoSe2-WSe2 heterostructure, composed of three WSe2 layers stacked atop three MoSe2 layers along the crystalline $c$-axis. Each monolayer possesses $D_{3h}$ symmetry, with an in-plane polarisation $\bf{P}$ arising from M–Se dipoles (where M = Mo or W). The 2H stacking sequence results in a $2\pi$ rotation of adjacent layers, aligning their polar axes anti-parallel. In this heterostructure, the polarisation $\bf{P}$ of MoSe2 layers slightly exceeds that of WSe2 layers, reflecting the greater electronegativity difference between Mo and Se compared to W and Se. However, the stronger atomic spin-orbit interaction (SOI) of W dominates the spin physics of the system. Coupled with $\bf{P}$, this SOI generates an out-of-plane spin-orbit field, $B_{\text{SO}}\propto(\bf{P}\times\bf{k})_{z}$, which induces a Zeeman-like spin splitting of energy bands. The effect is most pronounced at the K points of the Brillouin zone, as shown in Fig. 1(b) and (c), where the spin splitting in WSe2 ($\Delta E^{\text{W}}_{\text{SO}}$) is more than twice that in MoSe2 ($\Delta E^{\text{Mo}}_{\text{SO}}$).
@@
At the interface, the mismatch in spin-orbit coupling strengths gives rise to a hierarchical valley-spin polarisation. The MoSe2 valence bands are entirely contained within the $\Delta E^{\text{W}}_{\text{SO}}$ gap of WSe2, as shown in Fig. 1(f). This behaviour reflects the in-plane orbital nature of the K-point states, which are dominated by $d_{\parallel}={d_{xy},d_{x^{2}-y^{2}}}$ orbitals. These orbitals exhibit negligible out-of-plane coupling, suppressing interlayer hybridisation and preserving the monolayer-like electronic structure at K. By contrast, states at the $\Gamma$ point are dominated by $d_{z^{2}}$ orbitals, which extend into the van der Waals gap and exhibit substantial interlayer overlap. This leads to strong hybridisation at $\Gamma$, forming a broad bonding–anti-bonding energy window across the heterostructure, as depicted in Fig.1(g). While time-reversal symmetry suppresses spin splitting at $\Gamma$, the resulting band structure at this point is highly sensitive to interlayer coupling, in contrast to the robust monolayer-like dispersion retained at K. This dichotomy is evident in the full low-energy band structure of the 6-layer system (Fig.1(h)), where the K valleys retain their original characteristics while the $\Gamma$ valley forms a cascade of hybridised bands.
This stark contrast in orbital-dependent coupling is a hallmark of 2H-type transition metal dichalcogenides generally. The details of the $d_{\parallel}$ K-point valence bands are largely equivalent in monolayer and bulk 2H-TMD systems due to their minimal out-of-plane hopping potential Riley et al. (2015); Mak et al. (2010); Xiao et al. (2012), whereas $d_{z^{2}}$-derived $\Gamma$-point states exhibit substantial out-of-plane coupling to drive significant $k_{z}$ dispersion in bulk materials  Riley et al. (2014); Latzke et al. (2015); Kim et al. (2016); Alarab et al. (2023); Le Févre et al. (2024); Chang et al. (2014), thus underpinning the indirect-to-direct bandgap transition observed in few-layer systems Mak et al. (2010); Sun et al. (2016); Ernandes et al. (2021). Indeed, even isolated 3ML TMDs (Fig. 1(d-e)) have severely-altered energetics relative to their monolayer variants, but only around $\Gamma$.
@@
This disparity highlights the importance of $\Gamma$-derived orbitals in band engineering, where artificial heterostructures can exploit interlayer hybridisation to tune the out-of-plane ($\parallel k_{z}$) dispersion by controlling layer composition, ordering, and alignment. Indeed, in the above simple case of a 6-layer MoSe2-WSe2 heterostructure, though discretized, the bonding-antibonding formation of states at and around the $\Gamma$ point closely approximate a $k_{z}$-integrated bulk band structure of a bulk 2H-TMD system. This implies that, with only a few layers, the bulk band dispersion can be significantly restored and even controllably altered to exhibit a mixed characteristic between two or more materials. Access to an extra dimension within which electronic properties can be tailored and controlled is of great relevance for nano-quantum device applications, where miniaturisation plays a major role. More obvious approaches to bulk-like band engineering would rely on the assembly of numerous monolayers with fixed relative orientations, stacked sequentially and periodically to replicate bulk periodicity in momentum space through the curation of a well-defined unit cell. By bypassing this labor-intensive approach to instead employ much simpler, two-component heterostructures in the intermediate layer regime, the same essential three-dimensional electronic behavior can be accessed within an architecture that lends itself more naturally to fine tuning with e.g. layer thickness or twist angles.

### v2

# Dimensional band engineering in asymmetrical van der Waals heterostructures

@@
Abstract
@@
Van der Waals materials enable the construction of atomically sharp interfaces between compounds with distinct crystal and electronic properties. This is dramatically exploited in moiré systems, where a lattice mismatch or twist between monolayers generates an emergent in-plane periodicity, giving rise to electronic properties absent in the constituent materials. In contrast, vertical superlattices, formed by stacking dissimilar materials in the out-of-plane direction on the nanometer scale, have received far less attention despite their potential to realize analogous emergent phenomena in three dimensions. Through angle-resolved photoemission spectroscopy and density functional theory, we investigate six-to-eight-layer transition metal dichalcogenide (TMD) heterostructures constructed from pairs of stacked few-layer materials. Counterintuitively, we find that even these single superlattice units can host fully-delocalised bands, evidencing a robust coherent interlayer coupling across lattice-mismatched interfaces over extended spatial scales. We show how uncompensated semimetallic phases and energetically-mismatched topological surface states are readily and exclusively stabilized within such asymmetrical architectures. These findings establish two-component heterostructures in the intermediate layer-regime as platforms to invoke and control unprecedented combinations and instances of the diverse quantum phases native to many-layer TMDs.
I INTRODUCTION
@@
Transition metal dichalcogenides (TMDs) are popular components for functional devices due to the ease in which their vast array of intrinsic properties can be augmented Pham et al. (2022); Chhowalla et al. (2013); Ahn (2020); Ciarrocchi et al. (2022); Manzeli et al. (2017); Regan et al. (2022). This is facilitated through anisotropic bonding, enabling both the isolation of TMD flakes of arbitrary thicknesses and their subsequent recombination into stacks with atomically sharp interfaces without the need for epitaxial registry Lisi et al. (2021); Pei et al. (2022); Nunn et al. (2023); Stansbury et al. ; Jones et al. (2022). These procedures can both alter the electronic properties inherent to the component compounds, and induce otherwise-inaccessible correlated phenomena through the fine tuning of the laboratory-defined relative orientations, or ‘twist angle’ Devakul et al. (2021); Guo et al. (2023). Despite van der Waals bonding, however, the interlayer hopping along the $c$-axis in TMDs is significant, driving orbital-dependent bandwidths of up to several electron volts in bulk systems Clark et al. (2018, 2022); Riley et al. (2014); Watson et al. (2024). This is yet to be exploited in artificial heterostructures Wan et al. (2024), with current research largely rooted to homo- and heterobilayer systems, and thus to TMD properties in the 2D limit. This precludes, for example, the fine-tuning of band gaps through band quantisation Splendiani et al. (2010); Mak et al. (2010); Sun et al. (2016), the induction of $k_{z}$-mediated topological phases Bahramy et al. (2018), or the modification of charge density wave and magnetic transitions dependent on interlayer coupling Zhang et al. (2025); Watson et al. (2019); Feng et al. (2018).
Here, through nano-focused angle-resolved photoemission (nano-ARPES) and first principles calculations, we show how vertical 6-8 layer heterostructures, constructed simply by interfacing pairs of few-layer TMD flakes, possess emergent bands sensitive to the entire spatial extent of the hybrid system. Despite the absence of well-defined periodicity along the asymmetric $c$-axis, the fully delocalized bands mimic discretized three-dimensional states of ordered systems, but coexist with layer-locked states to produce an overall-spatially dependent electronic structure capable of supporting electronic phenomena usually found only in bulk compounds.
@@
We give special focus to $k_{z}$-mediated topological phase transitions, triggered by the interfacing of the few-layer TMD flakes, which produce pairs of distinct topologically non-trivial surfaces that are each defined largely by a single component material, unlike any naturally existing topological system. Similar phenomena should be expected for asymmetric heterostructures constructed from any pair of TMDs, and thus our findings unveil an enormous parameter space within which customizable electronic frameworks, mixing 2- and 3D properties from across the TMD family, can be produced and fine-tuned.
II METHODS
To construct the heterostructures characterised in this study, few-layer flakes of graphite, MoSe2, WSe2 and NbSe2 were exfoliated onto PDMS film and transferred onto n-type Si substrates. After each stacking stage, the samples were rinsed with diisoproylamine, isopropanol and ethanol to remove contaminants, and annealed in a constant-flow Ar furnace at 200-250 degrees. These procedures were performed in an Ar environment for samples containing NbSe2, with a hot plate used in place of a furnace. The component flakes comprising the heterostructures have thicknesses determined through comparisons of photoluminescence spectra and optical images to those of known references. An additional, in situ metric is provided by photoemission spectra obtained by probing the non-overlapping regions of each flake where available, wherein the number of distinct quantized $k_{z}$-sub-bands within the $d_{z^{2}}$-manifold at $\Gamma$ matches the number of monolayers along the c-axis Chang et al. (2014); Le Févre et al. (2024).
All ARPES data were collected using either $p$-polarised or unpolarised synchrotron light of energies between 60 and 144 eV at the MAESTRO beamline of the Advanced Light Source. Samples were cooled to cryogenic temperatures following a further 3+ hour UHV anneal at 270C. The base pressure of the setup was better than $\sim$$\times1$0-10 mbar. All heterostructures were pre-characterised with NanoESCA III at FMMA (ROR: 04z91ja70) Fli following a 3 hour UHV anneal at 270C, where an extraction voltage of 12kV and a HeI plasma discharge lamp was used for electronic structure measurements at room temperature. Electronic structure calculations were performed within Density Functional Theory using the Perdew-Burke-Ernzerhof exchange-correlation functional Perdew et al. (1996), as implemented in the VASP programme package Kresse and Furthmüller (1996); Kresse and Joubert (1999). Due to the close similarity between the lattice parameters of 2H-WSe2 and 2H-MoSe2, the experimentally reported bulk structure of 2H-WSe2 Schutte et al. (1987) used for both compounds in constructing (WSe2)n/(MoSe2)m heterostructures. A supercell containing $n=3$ and $m=3$ or $5$ layers was employed, with a vacuum spacing of 15 Å to prevent spurious interactions between periodic images. The plane-wave cut-off energy was set to 400 eV. The Brillouin zone was sampled using a $\mathbf{k}$-point mesh of $10\times 10\times 1$. Spin and orbital projections were obtained using Wannier functions Mostofi et al. (2008); Souza et al. (2001), downfolded from the DFT Hamiltonian, with the valence $d$ orbitals of the transition metal atoms and the Se $4p$ orbitals chosen as projection centres.
III RESULTS
III.1 Theoretical Framework
Figure 1: Chemical nature of an anisotropic transition metal heterostructure. (a) Schematic representation of a (WSe2)3/(MoSe2)3 heterostructure. Due to the $D_{3h}$ symmetry of each monolayer, a net dipole moment $\pm\mathbf{P}$ is induced, oriented perpendicular to the stacking axis, and oppositely aligned with respect to the adjacent layer. (b) and (c) Out-of-plane spin-projected electronic band structures for monolayers of WSe2 and MoSe2, respectively. The parameter $\Delta E_{\text{SO}}^{\text{W/Mo}}$ denotes the valley-dependent spin-orbit splitting at the K-point of the Brillouin zone. (d) and (e) Schematic energy diagrams illustrating the bonding and anti-bonding interactions at the K and $\Gamma$ points, respectively, with their orbital character predominantly influenced by the in-plane $d_{||}=\{{d_{xy},d_{x^{2}-y^{2}}}\}$ and out-of-plane $d_{z^{2}}$ orbitals of W/Mo. (f) Orbital-projected electronic structure of the (WSe2)3/(MoSe2)3 heterostructure, highlighting the distinct contributions from the constituent layers. The bands at $\Gamma$ are numbered from low energy to high. (g) and (h) Photoemission spectra collected from isolated 3ML thick flakes of WSe2 and MoSe2 (using 108 and 64 eV photons, respectively, and symmetrized about $\Gamma$). The bands at $\Gamma$ are numbered from low energy to high.
@@ III.1 Theoretical Framework
Figure 1 overviews the mechanism underpinning our findings. Fig.1(a) illustrates the lattice structure of a 6-layer MoSe2-WSe2 heterostructure, composed of three WSe2 layers stacked atop three MoSe2 layers along the crystalline $c$-axis. Each monolayer possesses $D_{3h}$ symmetry, with an in-plane polarisation $\bf{P}$ arising from M–Se dipoles (where M = Mo or W). The 2H stacking sequence results in a $\pi$ rotation of adjacent layers, aligning their polar axes anti-parallel. In this heterostructure, the polarisation $\bf{P}$ of MoSe2 layers slightly exceeds that of WSe2 layers, reflecting the greater electronegativity difference between Mo and Se compared to W and Se. However, the stronger atomic spin-orbit interaction (SOI) of W dominates the spin physics of the system. Coupled with $\bf{P}$, this SOI generates an out-of-plane spin-orbit field, $B_{\text{SO}}\propto(\bf{P}\times\bf{k})_{z}$, which induces a Zeeman-like spin splitting of energy bands. The effect is most pronounced at the K points of the Brillouin zone, as shown in Fig. 1(b) and (c), where the spin splitting in WSe2 ($\Delta E^{\text{W}}_{\text{SO}}$) is more than twice that in MoSe2 ($\Delta E^{\text{Mo}}_{\text{SO}}$).
@@
At the interface, the mismatch in spin-orbit coupling strengths gives rise to a hierarchical valley-spin polarisation. The MoSe2 valence bands are entirely contained within the $\Delta E^{\text{W}}_{\text{SO}}$ gap of WSe2, as shown in Fig. 1(d). This behaviour reflects the in-plane orbital nature of the K-point states, which are dominated by $d_{\parallel}={d_{xy},d_{x^{2}-y^{2}}}$ orbitals. These orbitals exhibit negligible out-of-plane coupling, suppressing interlayer hybridisation and preserving the monolayer-like electronic structure at K. By contrast, states at the $\Gamma$ point are dominated by $d_{z^{2}}$ orbitals, which extend into the van der Waals gap and exhibit substantial interlayer overlap. This leads to strong hybridisation at $\Gamma$, forming a broad bonding–anti-bonding energy window across the heterostructure, as depicted in Fig.1(e). While time-reversal symmetry suppresses spin splitting at $\Gamma$, the resulting band structure at this point is highly sensitive to interlayer coupling, in contrast to the robust monolayer-like dispersion retained at K. This dichotomy is evident in the full low-energy band structure of the 6-layer system (Fig.1(f)), where the K valleys retain their original characteristics while the $\Gamma$ valley forms a cascade of hybridised bands.
Figure 2: Multi-layer asymmetric heterostructures constructed from MoSe2 and WSe2 flakes. (a-d) Overview of a (WSe2)3/(MoSe2)5 heterostructure. (a) Optical image of the heterostructure region. The few-monolayer thick top and bottom layers are highlighted in red and blue respectively. The overlap region of interest (ROI) is highlighted in purple. (b) Band dispersions of the ROI along axes close to the $\Gamma-K$ direction for two stated photon energies. (c) Density functional theory based slab calculations, projected through the full heterostructure (left) and to the top layer (right) along the $\Gamma-K$ direction. (d) $k_{z}$ dispersion (hvv=60-144 eV) at $k_{\parallel}=0$. Dashed lines indicate the slices in (b), and solid lines are a tight-binding model (TBM) of a bulk (WSe2/MoSe2)n system. An inner potential of 16 eV is chosen to best match the periodicity of the calculated band structure. $k_{z}$ sub-bands of the $d_{z^{2}}$ valence band maxima are numbered from low binding energy to high. (e-h) Equivalent datasets for a (MoSe2)3/(WSe2)3 heterostructure.
This disparity highlights the importance of $\Gamma$-derived orbitals in band engineering, where artificial heterostructures can exploit interlayer hybridisation to tune the out-of-plane ($\parallel k_{z}$) dispersion by controlling layer composition, ordering, and alignment. Indeed, in the above simple case of a 6-layer MoSe2-WSe2 heterostructure, though discretized, the bonding-antibonding formation of states at and around the $\Gamma$ point closely approximate a $k_{z}$-integrated bulk band structure of a bulk 2H-TMD system. This implies that, with only a few layers, the bulk band dispersion can be significantly restored and even controllably altered to exhibit a mixed characteristic between two or more materials. To emphasize the significance of this, in Fig. 1(g-h), we compare the calculated electronic structure of the interfaced system to experimental band structures of individual free-standing 3ML WSe2 and MoSe2 flakes. In each case, the number of $d_{z^{2}}$-derived bands at $\Gamma$, which differ only by their discrete $k_{z}$ quantum number, is tripled relative to the respective monolayers as the three dimensional, $k_{z}$ dispersive, band continuum of the bulk limit begins to take form. The theoretical prediction in Fig. 1(f) demonstrates that, by simply combining these 3ML components, the evolution towards a 3D band continuum is entirely unhindered despite the interface between dissimilar compounds, and electrons at $\Gamma$ become fully delocalised through all six layers despite the enhanced spatial scales.
@@
More obvious approaches to bulk-like band engineering would rely on the assembly of numerous monolayers with fixed relative orientations, stacked sequentially and periodically to replicate bulk periodicity in momentum space through the curation of a well-defined unit cell, similar to that of the naturally occurring 4Hb polymorph of TaSe2 with alternating 1T/1H layers Ribak et al. . By bypassing this labor-intensive approach to instead employ much simpler, two-component heterostructures in the intermediate layer regime, the same essential three-dimensional electronic behavior can be accessed within an architecture that lends itself more naturally to fine tuning with e.g. layer thickness or twist angles. More generally, access to an extra dimension within which electronic properties can be tailored and controlled is of great relevance for nano-quantum device applications, where miniaturisation plays a major role.
III.2 Delocalised band formation in two-component stacks